OpenStudy (anonymous):

Find the three arithmetic means between 10 and 18.

OpenStudy (anonymous):

like average, mean and mode? IDK

OpenStudy (anonymous):

arithmetic mean

OpenStudy (anonymous):

U got many questions xD ____________________________ i will give u an example and u will conclude it (a,b,c) are a.m if b is a.m between (a) and (c) so b=(a+c)/2 or 2b=a+c _____ so what did u conclude ?

OpenStudy (anonymous):

yes i do.. 38 questions to be exact.

OpenStudy (anonymous):

good ,so can u get the a,means ?

OpenStudy (anonymous):

U need to learn it in case it came in your test or something Os wont be their ,Anyways here is the solution : ___________________ let the first A.m =b and the two numbers a ,c so b =(a+c)/2=(10+18)/2=14---->first a.m A.S(10,14,18) and soo on u can get the 2 rest A.meanings

OpenStudy (anonymous):

im not suree

OpenStudy (apoorvk):

okay what can one mean be? their average? the average of the two nos. 10 and 18?

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

so, what is the average of 10 and 18?

OpenStudy (anonymous):

8?

OpenStudy (apoorvk):

how do you guess that lol? average = (sum of nos.)/(no. of nos.)

OpenStudy (apoorvk):

here, sum of nos. = 10 + 18 = 28 no. of nos. = 2 so, avg. = 28/2 = 14 get this much?

OpenStudy (anonymous):

sort of?

OpenStudy (anonymous):

can you please just tell me wht it is?

OpenStudy (oaktree):

I'll tell you.

OpenStudy (oaktree):

12,14,16

OpenStudy (oaktree):

Look at it this way:

OpenStudy (oaktree):